OpenStudy (anonymous):

can someone help me figure out the next step. The equation is: 15/w = -10/w+5 so far I have 15(w+5) = -10w but am stuck as to what to do from here

OpenStudy (mertsj):

Use the distributive property

OpenStudy (anonymous):

i dont know how did you find the equation which is 15(w+5) = -10w but it is wrong. it can be solved in that way ; 15/w+10/w=5 25/w=5 w=5